# Foveated Retinotopy Improves Classification and Localization in CNNs

### v3

@@ Discussion
Fig 11: (A) An original image from the ImageNet dataset. (B) A heat map constructed by fitting bounding boxes to the image. (C) The heat map constructed in (B) and reduced to a resolution of $11\times 1$ 11 to be used as ground truth when evaluating the localisation performance of the networks.
ImageNet provides rich semantic links that allow the construction of task-specific datasets. It has previously been demonstrated that the use of fine-tuning to re-train networks such as Vgg16 [54] allows them to be applied to different tasks using the semantic network underlying ImageNet’s labels. Furthermore, it has been shown that the probability of a trained network performing a novel task (such as categorizing an animal) can be predicted using this semantic network, which links the outputs of a pre-trained network to a library of labels. This approach has been shown to be an effective method for learning to predict the presence of animals in images. [36]. We then exploit the semantic connection underlying the ImageNet labels by using these networks to perform a categorization task on the second dataset, the Animal 10k dataset, as referenced in Yu et al.’s work [55], with the goal of identifying the animals depicted in the images. The dataset provides a set of key points for each animal present in an image. For each image from Animal 10k containing a set of keypoints, we created a Gaussian heatmap centered on those points, with the peak value set to 1 and values decreasing with a standard deviation proportional to object size, thus capturing the true spatial extent and location of the target object within each image (see Figure 12). This approach allows for more effective localization and analysis of the visual distribution of animals in images.
Fig 12: (A) The original image of the Animal 10k dataset. (B) A heat map constructed by fitting Gaussians to the key points of the Animal 10k dataset. (C) The heat map constructed in (B) is normalized and reduced to a resolution of $11\times 1$ 11 to be used as ground truth when evaluating the heat map. A threshold (0.2) is applied to reduce the heat map field to the assumed contour of the animal.
@@
In this study, three series of transformations were used, depending on whether the network used the Cartesian or the retinotopic frame. In the case of the latter, a log-polar grid is used to facilitate the transformation of the image into the retinotopic frame. Due to the intrinsic nature of the transformation, which results in the cropping of a circular sample within the original image, a circular crop is implemented for the Cartesian frame to ensure comparability. To allow for a more straightforward comparison, the “raw” data sets were processed without a circular mask or polar logarithmic transformation. Each set underwent a uniform transformation, including normalization to tensors and resizing to a resolution of $244\times 244$ to match the pre-trained parameters of the model.
@@
In this study, we investigate the robustness of the deep learning models to natural image transformations that are easily perceived by humans. In particular, we evaluated the performance of the networks on the ImageNet dataset when the images were rotated by different angles and averaged the accuracy for each angle. To further evaluate the robustness to rotations, we also designed a “rotation-based attack”. To perform such an attack on a model $m$, we follow this simple procedure. Given an image $I$ and the output of the model ${\bm{p}}=m(I)$, which returns a probability vector over $K=1000$ classes, the loss function $\mathcal{L}$ is defined as the cross-entropy between the predicted probability vector and the ground truth label $y$, which we denote as $\mathcal{L}(m(I),y)$. This is the loss minimized during gradient descent training. We then systematically rotated the images and tracked the change in model loss. By denoting a rotation of the image by an angle $\theta$ as $\text{rot}(I;\theta)$, we define the rotation-based attack as the following heuristic for each image in the dataset:
@@
A widely accepted technique for evaluating the performance of Convolutional Neural Networks (CNNs) in localization tasks is the Class Activation Mapping (CAM) method. CAM works by analyzing the output of the CNN with respect to the target class, assigning weights to activations in each spatial feature map. This process generates a heat map that highlights significant areas of the image based on their contribution to the prediction. Building on the foundation of CAM, several derivative methods have emerged, including Grad-CAM [56], Score-CAM [57], and Opti-CAM [58].
@@
Fig 13: We illustrate the protocol used to obtain the likelihood map with an example on a single image (B). We show one sample for each sample ratio size ($1:10$ to $1:1$, from left to right). (A) The samples are cropped on the image in the Cartesian frame, we use circular cropping to match the area covered by the log-polar frame. (C) The corresponding samples cropped on the image in the retinotopic frame.
We used two parameters to define these maps. The first parameter is the resolution of the fixation point grid, and when not specified this resolution is set to $11\times 1$ 11. The second is the size of the samples clipped at each of these positions, defined as the ratio of the log-polar grid radius of the input to the total input size or Cartesian grid size, since the grid is a square for Cartesian samples (for an illustration of this simple process, see Appendix Figure 13). At each viewpoint, the largest possible sample is cropped. Thus a minimum sample with $1:10$ ratio at the border and the whole image at the center. From the Cartesian or retinotopic reference frame, this sample is then resized, if necessary, to a $224\times 224$ resolution to match the input size for the CNN before processing or transformed by the retinotopic mapping (also at $224\times 224$ resolution) before being used as input for the corresponding network, see Appendix Figure 13-B & D. Conveniently, a collection of samples for different fixation points can be processed as a single batch. This protocol defines a likelihood map for any given network as the likelihood of categorizing the presence of a label of interest inferred at regularly spaced fixation points in the image.

### v4

@@ Discussion
The dataset provides a set of key points for each animal present in an image. For each image from Animal 10k containing a set of keypoints, we created a Gaussian heatmap centered on those points, with the peak value set to 1 and values decreasing with a standard deviation proportional to object size, thus capturing the true spatial extent and location of the target animal within each image (see Figure 12). This approach allows for more effective localization and analysis of the visual distribution of animals in images.
Figure 12: (A) Original image of the Animal 10k dataset. (B) Heat map generated by fitting Gaussians to annotated key points. (C) Heat map from (B) normalized and downsampled to $11\times 11$, used as ground truth. A threshold of 0.2 is applied to delineate the assumed animal contour.
@@
In this study, three series of transformations were used, depending on whether the network used the Cartesian or the retinotopic reference frame. In the case of the latter, a log-polar grid is used to facilitate the transformation of the image into the retinotopic frame. Due to the intrinsic nature of the transformation, which results in the cropping of a circular sample within the original image, a circular crop is implemented for the Cartesian frame to ensure comparability. To allow for a more straightforward comparison, the “raw” datasets were processed without a circular mask or polar logarithmic transformation. Each set underwent a uniform transformation, including normalization to tensors and resizing to a resolution of $244\times 244$ to match the pre-trained parameters of the model.
@@
A common approach to evaluating the robustness of deep learning models is to subject them to adversarial attacks. In this study, we investigate the robustness of the deep learning models to natural image transformations that are easily perceived by humans. In particular, we evaluated the performance of the networks on the ImageNet dataset when the images were rotated by different angles and averaged the accuracy for each angle. To further evaluate the robustness to rotations, we also designed a “rotation-based attack” scenario. To perform such an attack on a model $m$, we follow this simple procedure. Given an image $I$ and the output of the model ${\bm{p}}=m(I)$, which returns a probability vector over $K=1000$ classes, the loss function $\mathcal{L}$ is defined as the cross-entropy between the predicted probability vector and the ground truth label $y$, which we denote as $\mathcal{L}(m(I),y)$. This is the loss minimized during gradient descent training. We then systematically rotated the images and tracked the change in model loss. By denoting a rotation of the image by an angle $\theta$ as $\text{rot}(I;\theta)$, we define the rotation-based attack as the following heuristic for each image in the dataset:
@@
A widely accepted technique for evaluating the performance of Convolutional Neural Networks (CNNs) in localization tasks is the Class Activation Mapping (CAM) method. CAM works by analyzing the output of the CNN with respect to the target class, assigning weights to activations in each spatial feature map. This process generates a heat map that highlights significant areas of the image based on their contribution to the prediction. Building on the foundation of CAM, several derivative methods have emerged, including Grad-CAM (Selvaraju et al., 2020), Score-CAM (Wang et al., 2020), and Opti-CAM (Zhang et al., 2024).
@@
Figure 13: We illustrate the protocol used to obtain the likelihood map with an example on a single image (B). We show one sample for each sample ratio size ($1:10$ to $1:1$, from left to right). (A) The samples are cropped on the image in the Cartesian frame, we use circular cropping to match the area covered by the log-polar frame. (C) The corresponding samples cropped on the image in the retinotopic frame.
@@
We used two parameters to define these maps. The first parameter is the resolution of the fixation point grid, and when not specified this resolution is set to $11\times 11$. The second is the size of the samples clipped at each of these positions, defined as the ratio of the log-polar grid radius of the input to the total input size or Cartesian grid size, since the grid is a square for Cartesian samples (for an illustration of this simple process, see Appendix Figure 13). At each viewpoint, the largest possible sample is cropped. Thus a minimum sample with $1:10$ ratio at the border and the whole image at the center. From the Cartesian or retinotopic reference frame, this sample is then resized, if necessary, to a $224\times 224$ resolution to match the input size for the CNN before processing or transformed by the retinotopic mapping (also at $224\times 224$ resolution) before being used as input for the corresponding network, see Appendix Figure 13-B & D. Conveniently, a collection of samples for different fixation points can be processed as a single batch. This protocol defines a likelihood map for any given network as the likelihood of categorizing the presence of a label of interest inferred at regularly spaced fixation points in the image.